OpenStudy (anonymous):

Eliminate the “z” variable in the first two equations by finding opposite coefficients and adding. What is the resulting equation? 4x – y + z = 6 3x + 2y – z = 0 4x – y – 3z = 2 a. 7x + y = 6 b.x – y = 6 c. x + y = 6 d.7x – y = 6

OpenStudy (anonymous):

4x – y + z = 6 3x + 2y – z = 0 7x-y=6

OpenStudy (anonymous):

4x - y + z = 6 3x + 2y - z = 0 so here we can see that the coefficient of the z terms are both one so we can apply addition straight away. 4x + 3x -y + 2y z - z 6 + 0 7x + y = 6 august I think u made a mistake in the y, it should be positive.

OpenStudy (anonymous):

so answer should be A.

OpenStudy (anonymous):

oh yes my bad